Given only a 30 Ohm and one 50 Ohm resister, list all possible values of resistance that can be obtained.
1
Expert's answer
2011-03-01T13:20:21-0500
You can connect them in two ways: in serial and parallel connection.
For serial: Rser =R1 + R2 = 30 + 50 = 80 [Ohm]
For parallel: Rpar = (R1R2)/(R1 + R2) = 18.75 [Ohm]
If only one resister is included into the chain, R1 = 30 [Ohm] and R2 = 50 [Ohm]